A can write 75 pages in 25 hours. A and B together can write 135 pages in 27 hours. in what time can B write 42 pages ?
Answer by TimothyLamb(4379)

rate = w/t
---
person a:
x = 75/25
x = 3 pg/hr
---
person b:
rate = y
---
x + y = 135/27
3 + y = 5
y = 2 pg/hr
---
t = w/rate
---
answer:
person b can write 42 pages in 42/2 = 21 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php